OpenStudy (anonymous):

derivative y = sqrt(3x+2)/(csc^2(x^3+1) solve with logarithmic differentiation

OpenStudy (kainui):

Show me your best attempt. Remember, square root is the same thing as to the 1/2 power.

OpenStudy (anonymous):

[(1/2\ln(3x+2)) / 2lncsc(x^3+1)] im not sure what to do after this

OpenStudy (anonymous):

There's a mistake with the log property you used.. \[\begin{align*}y&=\frac{\sqrt{3x+2}}{\csc^2(x^3+1)}\\ \ln y&=\ln\frac{\sqrt{3x+2}}{\csc^2(x^3+1)}\\ \ln y&=\ln(3x+2)^{1/2}-\ln\csc^2(x^3+1)\\ \ln y&=\frac{1}{2}\ln(3x+2)-2\ln\csc(x^3+1)\end{align*}\] Next step would be implicit differentiation.

OpenStudy (anonymous):

why did the division between the 3x+2 and the scs disappear?

OpenStudy (anonymous):

That's the property: log of a quotient is becomes the difference of logs. Mathematically, that's \[\ln\frac{a}{b}=\ln a-\ln b\]